OpenStudy (anonymous):

Convert 2.4 kg into cg. Please show your work.

OpenStudy (anonymous):

@cody7744

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

Convert 2.4 kg into cg. Please show your work. 100 cg = 1 g & 1 g = 0.001 kg 2.4kg = 2400g 2400 X 100 = 240000cg

OpenStudy (texaschic101):

1 kg /1000000 cg = 2.4 kg / x cg cross multiply x = 2.4(100,000) x = 240,000 2.4 kg = 240,000 cg

OpenStudy (anonymous):

Its the same thing...

OpenStudy (anonymous):

texaschic101 if she already answerd the question why do you need to reanswer it

OpenStudy (anonymous):

Go find another question to answer

OpenStudy (anonymous):

Thank y'all very much!

OpenStudy (texaschic101):

oops...put to many zeros 1/100000 = 2.4/x.....you will get the same answer

OpenStudy (anonymous):

Np glad to help

OpenStudy (texaschic101):

thats what we are here for :)

OpenStudy (texaschic101):

who ?

OpenStudy (anonymous):

@texaschic101 is just trying to help. I tagged her in it:)

OpenStudy (anonymous):

Can you pease go answer another question or you gettin reported

OpenStudy (anonymous):

Stop it @NessaCox

OpenStudy (texaschic101):

report me for what ?